Test TeXML

June 23, 1999

This is a section

This is a paragraph in the section.

This is a subsection

This is a paragraph in the subsection.

This is a subsubsection

This is a paragraph in the subsubsection.

Test descriptive list

mango a tropical fruit

honeydew a kind of melon

Test the escapes

Now we test all of the specials. The percent character, ’%’ is a comment character in TeX. The caret, ’ˆ ’ is a superscript. The amphersand, ’&’ is special for both XML and TeX. So are the angle braces, ’<’ and ’>’. The underscore, ’ ’ is a subscript. Dollar sign ’$’ puts TeX in math mode. Pound ’#’ signals a macro parameter. TeX interprets these as grouping characters, ’{’, ’}’. Finally, here’s good old active space, ’˜ ’ and the escape character, ’\’.